OpenStudy (anonymous):

Integral (2x)/(x-4) dx

OpenStudy (anonymous):

I tried letting u=x-4. And the x=u+4. Can this work?

OpenStudy (anonymous):

u= x-4 x=u+4 2x=2u +8 so the integral becomes integral of (2u+8)/u

OpenStudy (anonymous):

yes that will work

OpenStudy (anonymous):

Then I get integral 2 + (8/u) du

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

2u + 8Ln(u) +C